Let (x,d) be metric space and A proper subset of X .Define the closure of a set A .consider the usual metric space (Rn,d) .let A = {(x1,x2,.......xn): xi element of Q}


1
Expert's answer
2021-08-02T16:39:39-0400

There is three possible definitions of a closure of a set "A" (which are equivalent for a metric space) :

  • the smallest closed set which contains "A",
  • the set of all possible limits in "X" of sequences of elements of "A",
  • the set of points "x\in X" such that any neighbourhood of "x" intersects "A".

For a space "\mathbb{R}^n" with a usual metric "d" and a set "A" defined as "x\in A \leftrightarrow (x=(x_1,...,x_n) \wedge \forall i\: x_i \in \mathbb{Q})" the closure of "A" is the entire space "\mathbb{R}^n". We can prove it, using the second definition. Let "x\in\mathbb{R}^n" with "(x_1,..., x_n)" its components. By density of "\mathbb{Q}" in "\mathbb{R}", there are sequences "(x_i^k)_{k\in \mathbb{N}}" with "x^k_i \in \mathbb{Q}" and "x_i^k\overset{k\to+\infty}{\to} x_i" for all "1\leq i\leq n". The sequence in "A" defined as "((x_1^k,...,x_n^k))_{k\in\mathbb{N}}" converges to "x=(x_1,...,x_n)", as all coordinates of "x^k" converge to a respective coordinate of "x". Therefore, "x\in \bar A" and as this point is arbitrary, we have "\bar A=\mathbb{R}^n".